OpenStudy (anonymous):

Factor x^2 + 4x -21 (X+7)(x-3) -7)(+3 -7)(-3 +7)(-3

OpenStudy (anthonyn2121):

x^2+4x-2 factors into (x+7)(x-3). Set each binomial = to zero. x+7=0 x=-7. x-3=0 x=3